### v2

# Graded Betti numbers of the Jacobian algebra and total Tjurina numbers of plane curves

Abstract.
In this paper we compute an explicit closed formula for the total Tjurina number $\tau(C)$ of a reduced projective plane curve $C$ in terms of the graded Betti numbers of the corresponding Jacobian algebra. This formula allows a completely new view point on the classical upper bounds for the total Tjurina number $\tau(C)$ of a plane curve $C$ given by A. du Plessis and C. T. C. Wall. This approach yields in particular a new necessary condition for a set of positive integers to be the graded Betti numbers of the Jacobian algebra of a reduced plane curve.
Key words and phrases:
Jacobian ideal, Jacobian algebra, exponents, Tjurina numbers, graded Betti numbers
2010 Mathematics Subject Classification:
Primary 14H50; Secondary 14B05, 13D02, 32S22
1 partial support from the project “Singularities and Applications” - CF 132/31.07.2023 funded by the European Union - NextGenerationEU - through Romania’s National Recovery and Resilience Plan.
1. Introduction
Let $S=\mathbb{C}[x,y,z]$ be the polynomial ring in three variables $x,y,z$ with complex coefficients, and let $C:f=0$ be a reduced curve of degree $d\geq 3$ in the complex projective plane $\mathbb{P}^{2}$. We denote by $J_{f}$ the Jacobian ideal of $f$, i.e. the homogeneous ideal in $S$ spanned by the partial derivatives $f_{x},f_{y},f_{z}$ of $f$, and by $M(f)=S/J_{f}$ the corresponding graded quotient ring, called the Jacobian (or Milnor) algebra of $f$.
The study of the Jacobian module $M(f)$ of a reduced plane curve $C:f=0$ has become a central tool in understanding the geometry and singularities of $C$. Among the numerical invariants extracted from $M(f)$, the type
$t(C)=d_{1}+d_{2}+1-d\geq 0$
plays a particularly important role, where $d_{1}\leq d_{2}\leq\cdots\leq d_{m}$ are the exponents of $C$, see for instance [2, 13] and $d=\deg f$. We recall that $t(C)=0$ (resp. $t(C)=1$) if and only if the curve $C$ is free (resp. plus-one generated), see [1, 2]. After introducing the type of a plane curve, T. Abe, P. Pokora and the first author, have obtained a classification into two classes for curves with $t(C)=2$, see [2]. Later, we have established a classification into four classes for curves with $t(C)=3$, see [13].
In this paper we show that in general the type $t$ curves fall naturally into $2^{t-1}$ classes, see Section 2. In Section 3, we a closed formula for the corresponding Tjurina number in terms of the graded Betti numbers of the Jacobian algebra $M(f)$, see Theorem 3.1, which is the main result of this note. This formula is simplified by the use of the type $t=t(C)$.
As an example, we study these classes associated to partitions in the case of reduced plane curves of type $t(C)=4$ in Section 4. We show that exactly eight classes occur, and for each of them we show the non-emptiness by providing at least one example, see Example 4.2.
In Section 5 we look from a completely new view point at the classical upper bounds for the total Tjurina number $\tau(C)$ of a plane curve $C$ given by A. du Plessis and C.T.C. Wall in terms of $d$ and $d_{1}$. We also revisit the notions of nearly free curve, minimal plus-one generated curve and maximal Tjurina curve from this view point. The use of the type $t=t(C)$ of the curve $C$ makes the formulas much easier to state and also to use, see for instance the formulas in Corollary 5.2.
This approach yields in particular a new necessary condition for a set of positive integers to be the graded Betti numbers of the Jacobian algebra of a reduced plane curve, see Remark 5.3.
By analogy to the characterization of nearly free curves recalled in Corollary 5.4 (1), one may introduce the class of nearly maximal Tjurina curves and describe the corresponding graded Betti numbers, see Corollary 5.6.
All the computations in the examples in this note were done using the software CoCoA [3] and SINGULAR [4].
2. Graded Betti numbers and partitions
Consider an ordered partition
$\pi\ :\ t=\epsilon_{1}+\cdots+\epsilon_{s},$
where $t,\epsilon_{1},\cdots,\epsilon_{s}$ are strictly positive integers.
For integers $t\geq 1$ and $1\leq s\leq t$, we denote by $p(t,s)$ the number of such ordered partitions of $t$ with exactly $s$ parts. Since such partitions correspond bijectively to choices of s1s-1 points among the t1t-1 points $\{1,2,\ldots,t-1\}$ on the real line, one has
$p(t,s)=\binom{t-1}{s-1}$.
The total number of partitions of $t$ is therefore
$p(t)=\sum_{s=1}^{t}p(t,s)=\sum_{s=1}^{t}\binom{t-1}{s-1}=2^{t-1}.$
Consider now the general form of the minimal resolution of the Milnor algebra $M(f)$ of a reduced plane curve $C:f=0$ which is assumed to be not free:
(2.1) $0\longrightarrow\bigoplus_{j=1}^{m-2}S(-c_{j})\longrightarrow\bigoplus_{i=1}^{m}S(1-d-d_{i})\longrightarrow S^{3}(1-d)\longrightarrow S,$
with $c_{1}\leq...\leq c_{m-2}$ and $d_{1}\leq...\leq d_{m}$. It follows from [17, Lemma 1.1] that
(2.2) $c_{j}=d+d_{j+2}-1+\epsilon_{j},\qquad j=1,\dots,m-2,$
for some integers $\epsilon_{j}\geq 1$.
We call the ordered sequence of degrees $d_{i}$ for $i=1,\ldots,m$ and the integers $\epsilon_{j}$ for $j=1,\ldots,m-2$ for $m\geq 3$ the graded Betti numbers of the Jacobian algebra $M(f)$, since they determine and are determined by the usual graded Betti numbers of the Jacobian algebra $M(f)$ as defined for instance in [15].
Using [17, Formula (13)], one obtains the relation
(2.3) $d_{1}+d_{2}=d-1+\sum_{j=1}^{m-2}\epsilon_{j}.$
Hence we get an ordered partition
(2.4) $\pi_{C}\ :\ t(C)=\sum_{j=1}^{m-2}\epsilon_{j}.$
In other words, each curve $C$ with $t(C)=t$ induces an ordered partition $\pi_{C}$ of the integer $t\geq 1$. It is natural therefore to divide the curves of type $t$ into $2^{t-1}$ classes, each class, say denoted by $tT_{\pi}$, being determined by an ordered partition $\pi$ on $t$. More precisely, we say that $C$ has type $tT_{\pi}$ if $t(C)=t$ and $\pi_{C}=\pi$.
We put forth the following.
Conjecture 2.1.
For each $t\geq 1$ and each ordered partition $\pi$ of $t$, the set of plane curves of class $tT_{\pi}$ is not empty.
This conjecture holds for $t=1$ since in this case we have only one class and for $t=2$ (resp. $t=3$) as shown in [2] (resp. [13]). Example 4.2 below shows that this conjecture also holds for $t=4$. A much more difficult problem seems to be the following.
Question 2.2.
Find the necessary and sufficient conditions on the ordered sequence of degrees $d_{i}$ for $i=1,\ldots,m$ and on the integers $\epsilon_{j}\geq 1$ for $j=1,\ldots,m-2$, where $m\geq 3$, such that there is a reduced, non free curve $C:f=0$ with these data the graded Betti numbers of the Jacobian algebra $M(f)$.
Note that the degree of such a curve $C$ is determined by (2.3) and the inequality $c_{j}\leq c_{j+1}$ stated in (2.1) yields via (2.2) the necessary conditions
(2.5) $d_{j+2}+\epsilon_{j}\leq d_{j+3}+\epsilon_{j+1}$
for $j=1,\ldots,m-3$. We also know that
(2.6) $d_{1}\leq d_{2}\leq d_{3}\leq d-1,$
see for instance [9, Theorem 2.4] and
(2.7) $d_{m}\leq 2d-4,$
see for instance [6, inequality (2.6)]. In this note we prove the following inequality
(2.8) $E=d_{1}^{2}+td_{2}-\frac{t^{2}+\sum_{j=1}^{n}\epsilon_{j}^{2}}{2}-\sum_{j=1}^{n}\epsilon_{j}d_{j+2}\geq 0,$
which is a new necessary condition on these Betti numbers. In other words, (2.8) is not a consequence of the inequalities (2.5), (2.6) and (2.7), see for details Remark 5.3 below.
3. Tjurina numbers for curves via graded Betti numbers of their Jacobian algebra
Theorem 3.1.
Let $C:f=0$ be a reduced plane curve of degree $d$ and type $t(C)=t\geq 1$ with exponents $(d_{1},d_{2},\ldots,d_{m})$. Let $\pi_{C}=(\epsilon_{1},\dots,\epsilon_{n})$ be the corresponding ordered partition of $t$, with n=m2n=m-2. Then one has
$\tau(C)=d_{1}^{2}+d_{1}d_{2}+d_{2}^{2}-t(d_{1}+d_{2})-\sum_{i=3}^{m}\epsilon_{i-2}d_{i}+R,$
where
$R=\frac{t^{2}-\sum_{i=3}^{m}\epsilon_{i-2}^{2}}{2}=\sum_{3\leq i<j\leq m}\epsilon_{i-2}\epsilon_{j-2}\geq 0.$
In particular, $R=0$ if and only if $m=3$.
Proof.
By taking the homogeneous component of degree $k$ and using exactness in the resolution (2.1) we get the following equality, for $k\gg 0$.
$\dim M(f)_{k}=\binom{k+2}{2}-3\binom{k+3-d}{2}+\sum_{i=1}^{n+2}\binom{k+3-d-d_{i}}{2}-$
$\sum_{j=1}^{n}\binom{k+3-d-d_{j+2}-\epsilon_{j}}{2}$.
This is a polynomial in $k$ of degree at most $2$. The constant term of this polynomial is $\tau(C)$, and a direct computation, done using $d=d_{1}+d_{2}+1-t$, the equality (2.4), and grouping the terms in the above sum as follows
$\dim M(f)_{k}=\binom{k+2}{2}-3\binom{k+3-d}{2}+\sum_{i=1}^{2}\binom{k+3-d-d_{i}}{2}$+
$\sum_{j=1}^{n}\left(\binom{k+3-d-d_{j+2}}{2}-\binom{k+3-d-d_{j+2}-\epsilon_{j}}{2}\right)$
yields the result.
Remark 3.2.
(i) Theorem 3.1 is a generalization of [2, Proposition 1.11] which treats the case $t=2$ and of [13, Theorem 3.1] which treats the case $t=3$.
(ii) When $C$ is a free curve of degree $d$ with exponents $(d_{1},d_{2})$, then $d_{1}+d_{2}=d-1$ and
$\tau(C)=(d-1)^{2}-d_{1}(d-d_{1}-1)=(d_{1}+d_{2})^{2}-d_{1}d_{2}=d_{1}^{2}+d_{1}d_{2}+d_{2}^{2},$
see for instance [7]. Since in this case $t(C)=0$, we see that the equality in Theorem 3.1 holds in this case as well with the obvious conventions
$\sum_{i=3}^{m}\epsilon_{i-2}d_{i}=\sum_{i=3}^{m}\epsilon_{i-2}^{2}=0.$
Indeed, in this case there are no $\epsilon_{j}$ since n=m2=0n=m-2=0.
Example 3.3.
(i) For the partition $\pi=(t)$ of $t(C)=t$, we get the formula
$\tau(C)=d_{1}^{2}+d_{1}d_{2}+d_{2}^{2}-t(d_{1}+d_{2}+d_{3}).$
Indeed, in this case $m=3$ and $\epsilon_{1}=t$. To have a geometric example, if $C$ is a smooth curve, then the exponents of $C$ are
(d1,d1,d1)(d-1,d-1,d-1)
and the type of $C$ is t(C)=d1t(C)=d-1. If we consider the partition $\pi=(d-1)$, then clearly the smooth curve $C$ has class $(d-1)T_{\pi}$.
(ii) For the partition $\pi^{\prime}=(1,1,\ldots,1)$ of $t(C)=t$, we get the formula
$\tau(C)=d_{1}^{2}+d_{1}d_{2}+d_{2}^{2}-t(d_{1}+d_{2})-\sum_{j=3}^{t+2}d_{j}+\frac{t(t-1)}{2}.$
In this case $m=t+2$ and $\epsilon_{1}=\ldots=\epsilon_{t}=1$. Note that the maximal Tjurina curves discussed in Corollary 5.5 below are in this class of curves.
4. Examples for curves of type $t(C)=4$
The following is a direct consequence of Theorem 3.1. Consider all the possible partitions of $t$, namely:
$\pi_{1}=(4),\ \pi_{2}=(3,1),\ \pi_{3}=(2,2),\ \pi_{4}=(1,3),\ \pi_{5}=(2,1,1),\ \pi_{6}=(1,2,1),$
$\pi_{7}=(1,1,2)\text{ and }\pi_{8}=(1,1,1,1).$
Corollary 4.1.
Let $C$ be a reduced plane curve of type $t(C)=4$ with exponents $(d_{1},d_{2},\ldots,d_{m})$. Then $d_{1}+d_{2}=d+3$, and $C$ is in exactly one of the following eight classes.
  1. Type $4T_{\pi_{1}}$.
    $C$ is a $3$-syzygy curve and $\epsilon_{1}=4$. If $(d_{1},d_{2},d_{3})$ are the exponents of $C$, then
    $\tau(C)=d_{1}^{2}+d_{1}d_{2}+d_{2}^{2}-4(d_{1}+d_{2}+d_{3}).$
  2. Type $4T_{\pi_{2}}$.
    $C$ is a $4$-syzygy curve and $\epsilon_{1}=3$, $\epsilon_{2}=1$. If $(d_{1},d_{2},d_{3},d_{4})$ are the exponents of $C$, then $d_{3}<d_{4}-1$ and
    $\tau(C)=d_{1}^{2}+d_{1}d_{2}+d_{2}^{2}-4d_{1}-4d_{2}+3-3d_{3}-d_{4}.$
  3. Type $4T_{\pi_{3}}$.
    $C$ is a $4$-syzygy curve and $\epsilon_{1}=2$, $\epsilon_{2}=2$. If $(d_{1},d_{2},d_{3},d_{4})$ are the exponents of $C$, then
    $\tau(C)=d_{1}^{2}+d_{1}d_{2}+d_{2}^{2}-4d_{1}-4d_{2}+4-2d_{3}-2d_{4}.$
  4. Type $4T_{\pi_{4}}$.
    $C$ is a $4$-syzygy curve and $\epsilon_{1}=1$, $\epsilon_{2}=3$. If $(d_{1},d_{2},d_{3},d_{4})$ are the exponents of $C$, then
    $\tau(C)=d_{1}^{2}+d_{1}d_{2}+d_{2}^{2}-4d_{1}-4d_{2}+3-d_{3}-3d_{4}.$
  5. Type $4T_{\pi_{5}}$.
    $C$ is a $5$-syzygy curve and $\epsilon_{1}=2$, $\epsilon_{2}=1$, $\epsilon_{3}=1$. If $(d_{1},d_{2},d_{3},d_{4},d_{5})$ are the exponents of $C$, then $d_{3}<d_{4}$ and
    $\tau(C)=d_{1}^{2}+d_{1}d_{2}+d_{2}^{2}-4d_{1}-4d_{2}+5-2d_{3}-d_{4}-d_{5}.$
  6. Type $4T_{\pi_{6}}$.
    $C$ is a $5$-syzygy curve and $\epsilon_{1}=1$, $\epsilon_{2}=2$, $\epsilon_{3}=1$. If $(d_{1},d_{2},d_{3},d_{4},d_{5})$ are the exponents of $C$, then $d_{4}<d_{5}$ and
    $\tau(C)=d_{1}^{2}+d_{1}d_{2}+d_{2}^{2}-4d_{1}-4d_{2}+5-d_{3}-2d_{4}-d_{5}.$
  7. Type $4T_{\pi_{7}}$.
    $C$ is a $5$-syzygy curve and $\epsilon_{1}=1$, $\epsilon_{2}=1$, $\epsilon_{3}=2$. If $(d_{1},d_{2},d_{3},d_{4},d_{5})$ are the exponents of $C$, then
    $\tau(C)=d_{1}^{2}+d_{1}d_{2}+d_{2}^{2}-4d_{1}-4d_{2}+5-d_{3}-d_{4}-2d_{5}.$
  8. Type $4T_{\pi_{8}}$.
    $C$ is a $6$-syzygy curve and $\epsilon_{1}=\epsilon_{2}=\epsilon_{3}=\epsilon_{4}=1$. If $(d_{1},d_{2},d_{3},d_{4},d_{5},d_{6})$ are the exponents of $C$, then
    $\tau(C)=d_{1}^{2}+d_{1}d_{2}+d_{2}^{2}-4d_{1}-4d_{2}+6-(d_{3}+d_{4}+d_{5}+d_{6}).$
Proof.
The only points that need additional explanations are the following. The inequality $d_{3}<d_{4}-1$ in Type $4T_{\pi_{2}}$ and the inequality $d_{3}<d_{4}$ in Type $4T_{\pi_{5}}$ come from the inequality stated in (2.5) for $j=1$ The inequality $d_{4}<d_{5}$ in Type $4T_{\pi_{6}}$ comes from the similar inequality stated in (2.5) for $j=2$. ∎
We give next explicit examples of reduced plane curves of type $t(C)=4$, one for each of the eight possible classes listed above. For each example we list the degree $d$, the Tjurina number $\tau(C)$, the exponents of $C$, and the minimal free resolution of the Milnor algebra $M(f)$.
Example 4.2.
(i) Type $4T_{\pi_{1}}$ Consider the curve
$C:f=x^{5}+y^{5}+z^{5}=0.$
Then $d=5$, $t(C)=4$, $m=3$, $\tau(C)=0$, and the exponents are $(4,4,4)$. The minimal free resolution of $M(f)$ is
$0\to S(-12)\to S(-8)^{3}\to S(-4)^{3}\to S.$
(ii) Type $4T_{\pi_{2}}$ Consider the curve
$C:f=x^{5}+xyz^{3}+z^{5}=0.$
Then $d=5$, $t(C)=4$, $m=4$, $\tau(C)=1$, and the exponents are $(4,4,4,6)$. The minimal free resolution of $M(f)$ is
$0\to S(-11)^{2}\to S(-8)^{3}\oplus S(-10)\to S(-4)^{3}\to S.$
(iii) Type $4T_{\pi_{3}}$
Consider the curve
$C:f=(x^{3}+y^{3})^{3}+(y^{3}+z^{3})^{3}=0.$
Then $d=9$, $t(C)=4$, $m=4$, $\tau(C)=36$, and the exponents are $(4,8,8,8)$. The minimal free resolution of $M(f)$ is
$0\to S(-18)^{2}\to S(-12)\oplus S(-16)^{3}\to S(-8)^{3}\to S.$
(iv) Type $4T_{\pi_{4}}$
Consider the curve
$C:f=x^{6}+y^{4}z^{2}+z^{4}x^{2}=0.$
Then $d=6$, $t(C)=4$, $m=4$, $\tau(C)=8$, and the exponents are $(4,5,5,5)$. The minimal free resolution of $M(f)$ is
$0\to S(-11)\oplus S(-13)\to S(-9)\oplus S(-10)^{3}\to S(-5)^{3}\to S.$
(v) Type $4T_{\pi_{5}}$ Consider the curve
$C:f=x^{7}+y^{7}-4x^{2}y^{2}z^{3}=0.$
Then $d=7$, $t(C)=4$, $m=5$, $\tau(C)=14$, and the exponents are $(5,5,6,7,7)$. The minimal free resolution of $M(f)$ is
$0\to S(-14)^{3}\to S(-11)^{2}\oplus S(-12)\oplus S(-13)^{2}\to S(-6)^{3}\to S.$
(vi) Type $4T_{\pi_{6}}$ Consider the curve
$C:f=z(x+y-z)(2x+3y-5z)(3x+5y-7z)(x^{3}+y^{3}-z^{3})=0.$
Then $d=7$, $t(C)=4$, $m=5$, $\tau(C)=19$, and the exponents are $(5,5,5,5,6)$. The minimal free resolution of $M(f)$ is
$0\to S(-12)\oplus S(-13)^{2}\to S(-11)^{4}\oplus S(-12)\to S(-6)^{3}\to S.$
(vii) Type $4T_{\pi_{7}}$ Consider the curve
$C:f=(x^{4}+y^{4})^{2}-4x^{2}y^{2}z^{4}=0.$
Then $d=8$, $t(C)=4$, $m=5$, $\tau(C)=28$, and the exponents are $(5,6,6,6,6)$. The minimal free resolution of $M(f)$ is
$0\to S(-14)^{2}\oplus S(-15)\to S(-12)\oplus S(-13)^{4}\to S(-7)^{3}\to S.$
(viii) Type $4T_{\pi_{8}}$ Consider the curve
$C:f=x^{3}y^{2}+y^{3}z^{2}+x^{2}z^{3}=0.$
Then $d=5$, $t(C)=4$, $m=6$, $\tau(C)=6$, and the exponents are $(4,4,4,4,4,4)$. The minimal free resolution of $M(f)$ is
$0\to S(-9)^{4}\to S(-8)^{6}\to S(-4)^{3}\to S.$
5. On a result by A. du Plessis and C.T.C. Wall
We recall first the following result due to du Plessis and Wall, see [14, Theorem 3.2] as well as [16] for an alternative approach.
Theorem 5.1.
For positive integers $d$ and $d_{1}$, define two new integers by
$\tau(d,d_{1})_{min}=(d-1)(d-d_{1}-1)\text{ and }\tau(d,d_{1})_{max}=(d-1)^{2}-d_{1}(d-d_{1}-1).$
If $C:f=0$ is a reduced curve of degree $d$ in $\mathbb{P}^{2}$ and $d_{1}$ is the first exponent of $f$, then
$\tau(d,d_{1})_{min}\leq\tau(C)\leq\tau(d,d_{1})_{max}$
and the equality $\tau(C)=\tau(d,d_{1})_{max}$ holds if and only if $2d_{1}<d$ and $C$ is a free curve. Moreover, for $2d_{1}\geq d$, the stronger inequality $\tau(C)\leq\tau^{\prime}(d,d_{1})_{max}$ holds, where
$\tau(d,d_{1})^{\prime}_{max}=\tau(d,d_{1})_{max}-\binom{2d_{1}+2-d}{2}$.
We recall that a maximal Tjurina curve $C$ is a curve such that $2d_{1}\geq d$ and $\tau(C)=\tau(d,d_{1})^{\prime}_{max}$, where see [10]. As an example, the curve in Example 4.2 (viii) is a maximal Tjurina curve with $d=5$ and $d_{1}=4$. Theorem 3.1 yields a precise way to measure the difference between $\tau(C)$ and $\tau(d,d_{1})_{max}$ (resp. $\tau(d,d_{1})^{\prime}_{max}$), and in this way gives, in particular, a proof of the claims involving these invariants in Theorem 5.1.
Corollary 5.2.
Let $C:f=0$ be a reduced plane curve of degree $d$ and type $t(C)=t\geq 1$ with exponents $(d_{1},d_{2},\ldots,d_{m})$ and $\pi_{C}=(\epsilon_{1},\dots,\epsilon_{n})$ the corresponding ordered partition of $t$, with n=m2n=m-2. Then one has the following.
  1. (1)
    For any curve $C$ as above, one has
    $\tau(d,d_{1})_{max}-\tau(C)=\frac{t^{2}+t}{2}+\frac{\sum_{j=1}^{n}(\epsilon_{j}^{2}-\epsilon_{j})}{2}+\sum_{j=1}^{n}\epsilon_{j}(d_{j+2}-d_{2}).$
  2. (2)
    If in addition $2d_{1}\geq d$, then
    $\tau(d,d_{1})^{\prime}_{max}-\tau(C)=\frac{t^{2}-(t-(d_{2}-d_{1}))^{2}}{2}+\frac{\sum_{j=1}^{n}(\epsilon_{j}^{2}-\epsilon_{j})+(d_{2}-d_{1})}{2}+\sum_{j=1}^{n}\epsilon_{j}(d_{j+2}-d_{2}).$
Moreover, all the terms in the right hand sides in (1) and (2) are positive.
Proof.
We give the proof only in the case (2), the case (1) being similar and simpler. Since $d_{1}+d_{2}\geq 2d_{1}>d-1$, it follows that $t=t(C)\geq 1$. Using the formula for $\tau(C)$ given in Theorem 3.1, the equality (2.4) and replacing $2d_{1}+2-d$ in the binomial coefficient by $t-(d_{2}-d_{1})+1$, we get
$(d-1)^{2}-d_{1}(d-d_{1}-1)-\binom{t-(d_{2}-d_{1})+1}{2}-\tau(C)=$
$\frac{t^{2}-(t-(d_{2}-d_{1}))^{2}}{2}+\frac{1}{2}\sum_{j=1}^{n}(\epsilon_{j}^{2}-\epsilon_{j})+\frac{d_{2}-d_{1}}{2}+\sum_{j=1}^{n}\epsilon_{j}(d_{j+2}-d_{2}).$
The right hand side is a sum where all the terms are positive. Indeed, note that
$t^{2}-(t-(d_{2}-d_{1})^{2}=(d_{2}-d_{1})(2t-(d_{2}-d_{1})=2(d_{2}-d_{1})(2d_{1}+1-d)\geq 0.$
Hence this sum is equal to 0 if and only if all the terms are equal to 0. This implies that
$d_{1}=d_{2}=\ldots=d_{m}$
and $\epsilon_{j}=1$ for all $j^{\prime}s$. Finally, the equality (2.3) yields $m=2d_{1}-d+3$, which completes our proof. ∎
Remark 5.3.
A direct computation using the formula for $\tau(C)$ given in Theorem 3.1 yields
$E=\tau(C)-\tau(d,d_{1})_{min}=d_{1}^{2}+td_{2}-\frac{t^{2}+\sum_{j=1}^{n}\epsilon_{j}^{2}}{2}-\sum_{j=1}^{n}\epsilon_{j}d_{j+2}.$
For $m=3$ the inequality $E\geq 0$ holds for obvious reasons, since one has $\epsilon_{1}=t$ and hence
$E=d_{1}(d_{1}-t)+t(d_{1}+d_{2}-t-d_{3})=d_{1}(d_{1}-t)+t(d-1-d_{3})\geq 0.$
Indeed, $d_{1}-t=d-1-d_{2}\geq 0$ and $d-1-d_{3}\geq 0$ by (2.6).
On the other hand, for $m\geq 4$, the inequality $E\geq 0$ coming from Theorem 5.1 is a new necessary condition to be satisfied by the exponents $(d_{1},d_{2},\ldots,d_{m})$ and $(\epsilon_{1},\dots,\epsilon_{n})$, the corresponding ordered partition of $t$.
To show that this inequality is not a consequence of the inequalities (2.5), (2.6) and (2.7), one may consider the data $d_{1}=3$, $d_{2}=d_{3}=8$, $d_{4}=15$, $\epsilon_{1}=\epsilon_{2}=1$ which verify the inequalities (2.5), (2.6) and (2.7), but not the above inequality. Indeed, for these data one gets E=1<0E=-1<0, which shows that there is no reduced curve $C$ with these Betti numbers.
Recall that an $3$-syzygy curve $C$ is called nearly free (resp. minimal plus-one generated) if $d_{1}+d_{2}=d$ and $d_{2}=d_{3}$ (resp. $d_{1}+d_{2}=d$ and $d_{2}+1=d_{3}$, see [8, 11]. Using Corollary 5.2 (1) we can give new proofs for the following known results. Indeed, the claim (1) below occurs in [5, Theorem 1.3], and the claim (2) below occurs in [11, Theorem 1.5].
Corollary 5.4.
Let $C:f=0$ be a reduced plane curve of degree $d$ with exponents $(d_{1},d_{2},\ldots,d_{m})$. Then one has the following.
  1. (1)
    $C$ is a nearly free curve if and only if $\tau(C)=\tau(d,d_{1})_{max}-1$.
  2. (2)
    $C$ is a minimal plus-one generated curve if and only if $\tau(C)=\tau(d,d_{1})_{max}-2$.
Proof.
If the curve $C$ is nearly free or respectively minimal plus-one generated, it follows from the general formula for the Tjurina number of a 3-syzygy curve given in [9] that $\tau(C)$ satisfies the claimed relations.
Conversely, if $\tau(C)$ satisfies the claimed relations, it follows from Theorem 5.1 that the curve $C$ is not free, and hence $t(C)\geq 1$.
It is clear that
$\tau(d,d_{1})_{max}-\tau(C)=1$
implies $t=1$, and hence $C$ is a plus-one generated curve, in particular $m=3$. Then we get $\epsilon_{1}=1$ and $d_{2}=d_{3}$, which tells us that $C$ is a nearly free curve.
Next, if we have
$\tau(d,d_{1})_{max}-\tau(C)=2,$
then again $t=1$ and $m=3$ as above. Then we get either $\epsilon_{1}=1$ or $\epsilon_{1}=2$. In the first case we get $d_{3}=d_{2}+1$, that is $C$ is a minimal plus-one generated curve. In the second case, we get $d_{2}=d_{3}$, that is $C$ is a nearly free curve. But this is impossible, since it contradicts (1).
Using Corollary 5.2 (2) we can give a completely new proof of [10, Theorem 3.1], that characterizes the maximal Tjurina curves in terms of their resolutions.
Corollary 5.5.
Let $C:f=0$ be a reduced plane curve of degree $d$ and type $t(C)=t\geq 1$ with exponents $(d_{1},d_{2},\ldots,d_{m})$ and $\pi_{C}=(\epsilon_{1},\dots,\epsilon_{n})$ the corresponding ordered partition of $t$, with n=m2n=m-2. Then $C:f=0$ is a maximal Tjurina curve if and only if the exponents of $C$ satisfy
$d_{1}=d_{2}=\ldots=d_{m},$
with $m=2d_{1}-d+3$ and
$\epsilon_{1}=\ldots=\epsilon_{n}=1.$
Proof.
The sum in Corollary 5.2 (2) is equal to 0 if and only if all the terms are equal to 0. This implies that
$d_{1}=d_{2}=\ldots=d_{m}$
and $\epsilon_{j}=1$ for all $j^{\prime}s$. Finally, the equality (2.3) yields $m=2d_{1}-d+3$, which completes our proof. ∎
We have the following analog of Corollary 5.4 (1), where $\tau(d,d_{1})_{max}$ is replaced by $\tau(d,d_{1})^{\prime}_{max}$. We say that $C:f=0$ is a nearly maximal Tjurina curve if $2d_{1}\geq d$ and
$\tau(C)=\tau(d,d_{1})^{\prime}_{max}-1.$
Corollary 5.6.
Let $C:f=0$ be a reduced plane curve of degree $d$ and type $t(C)=t\geq 1$ with exponents $(d_{1},d_{2},\ldots,d_{m})$ and $\pi_{C}=(\epsilon_{1},\dots,\epsilon_{n})$ the corresponding ordered partition of $t$, with n=m2n=m-2. Then $C$ is a nearly maximal Tjurina curve if and only if one of the following two situations occurs.
  1. (1)
    $d_{1}=d_{2}=\ldots=d_{m-1}$, $d_{m}=d_{1}+1$ and $\epsilon_{1}=\ldots=\epsilon_{n}=1$, with $m=2d_{1}+3-d$.
  2. (2)
    $d_{1}=d_{2}=\ldots=d_{m}$, $\epsilon_{1}=\ldots=\epsilon_{n-1}=1$ and $\epsilon_{n}=2$, with $m=2d_{1}+2-d$.
Proof.
Note that $t(C)\geq 1$ implies $m\geq 3$, which yields $2d_{1}\geq d$ in both cases (1) and (2). If one of the situations (1) or (2) occurs, then using Corollary 5.2 (2) it follows that $C$ is a nearly maximal Tjurina curve.
Conversely, assume now that $C$ is a nearly maximal Tjurina curve. Assume first that $d_{1}=d_{2}$. Then either all the $\epsilon_{j}$ are $1$, and this leads to the case (1), or all the $\epsilon_{j}$ are $1$ except one, which is $2$. Then we have $d_{1}=d_{2}=\ldots=d_{m}$ and we must have $\epsilon_{n}=2$ due to the inequality $c_{n-1}\leq c_{n}$ in (2.1). Hence this case corresponds to the claim (2) in our result.
Assume now that $d_{1}<d_{2}$. Then
$t^{2}-(t-(d_{2}-d_{1})^{2}=2(d_{2}-d_{1})(2d_{1}+1-d)\geq 2,$
as we have seen in the proof of Corollary 5.2. It follows that
$\frac{t^{2}-(t-(d_{2}-d_{1}))^{2}}{2}+\frac{d_{2}-d_{1}}{2}>1$
which is a contradiction. Hence this case cannot occur. ∎
Remark 5.7.
To state the conditions on the partition $\pi_{C}=(\epsilon_{1},\dots,\epsilon_{n})$ that occurs in Corollary 5.6, it is convenient to introduce the invariant of $C$ given by
$\Delta m=t-(m-2)=d_{1}+d_{2}-d-m+3,$
which was already considered in [12] when $d_{1}=d_{2}$. Using (2.4), it follows that
$\Delta m=\sum_{j=1}^{m-2}(\epsilon_{j}-1)\geq 0.$
It follows that the condition on the partition $\pi_{C}$ in Corollary 5.6 (1) is equivalent to
$\Delta m=0,$
and the condition on the partition $\pi_{C}$ in Corollary 5.6 (2) is equivalent to
$\Delta m=1,$
when the additional condition $d_{1}=d_{2}=\ldots=d_{m}$ holds. With the terminology introduced in [12] a curve $C$ satisfying $\pi_{C}$ in Corollary 5.6 (2) is exactly a curve $C$ of type $(d,d_{1},m)$ such that $\Delta m=1$.
Example 5.8.
Here are some examples of curves satisfying the conditions in Corollary 5.6.
(i) Consider the curve
C:f=xyz(xz)(x2z)(x3z)(yz)(y2z)(y3z)(x+y)(x+y2z)=0C:f=xyz(x-z)(x-2z)(x-3z)(y-z)(y-2z)(y-3z)(x+y)(x+y-2z)=0
from [12, Remark 4.3]. Then $C$ is a 4-syzygy curve with exponents $(6,6,6,7)$ and degree $d=11$. It follows that
$\Delta m=d_{1}+d_{2}-d-m+3=6+6-11-4+3=0,$
therefore by our Remark 5.7 this curve satisfies the conditions in Corollary 5.6 (1).
Similarly, the curve
$C^{\prime}:f^{\prime}=xyz(x+y)(y+z)(z+x)(-x+y+z)(x-y+z)(x+y-z)=0$
has exponents $(5,5,5,6)$ and $d=9$, hence again
$\Delta m=d_{1}+d_{2}-d-m+3=5+5-9-4+3=0.$
The irreducible quartic
$C^{\prime\prime}:f^{\prime\prime}=x^{4}+x^{3}z+y^{2}z^{2}=0$
has exponents $(2,2,3)$ and hence it is minimal plus-one generated as in Corollary 5.4. On the other hand it verifies
$\Delta m==d_{1}+d_{2}-d-m+3=2+2-4-3+3=0.$
(ii) Using the final part of Remark 5.7 we have the following. The curves $C$ and $C^{\prime}$ in [12, Example 4.5] satisfy the conditions in Corollary 5.6 (2) for $d=7$, $d_{1}=4$ and $m=3$. The curves $C$ in [12, Example 4.6] satisfies the conditions in Corollary 5.6 (2) for $d=9$, $d_{1}=5$ and $m=3$. The curves $C$ in [12, Example 4.7] satisfies the conditions in Corollary 5.6 (2) for $d=11$, $d_{1}=6$ and $m=3$. To get examples of curves satisfying the conditions in Corollary 5.6 (2) with $m\geq 4$, one may consider the curves $C$, $^{\prime}, C′′′C^{\prime\prime\prime}$ and $C_{6}$ in [12, Example 5.1], the curves $C_{7}$ and $C^{\prime}$ in [12, Example 5.2], the curves $C$ and $C_{8}$ in [12, Example 5.3] and the curves $C^{\prime}$ and $C^{\prime\prime\prime}$ in [12, Example 5.4].